OpenStudy (anonymous):

derivative of (x^(2))(e^(1/x))

OpenStudy (anonymous):

@Callisto

OpenStudy (anonymous):

i got (2x)(e^(1/x))+x^(2)(e^(1/x) which is wrong

OpenStudy (anonymous):

yeah it is

OpenStudy (anonymous):

\[g(x)=e^{\frac{1}{x}}\] \[g'(x)=-\frac{1}{x^2}e^{\frac{1}{x}}\] by the chain rule

OpenStudy (anonymous):

so the second part of your product rule is the problem. first part is good

OpenStudy (anonymous):

but how is it (-1/x^2) e^(1/x)?

OpenStudy (anonymous):

shouldn't the derivative e^(1/x)

OpenStudy (callisto):

\[y=e^{\frac{1}{x}}\] Let u = 1/x \[\frac{dy}{dx} = \frac{dy}{du}\times \frac{du}{dx} = \frac{d}{du}e^u \times \frac{d}{dx}(\frac{1}{x})= ...?\]

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

or 0

OpenStudy (anonymous):

by quotient rule i did (1)'(x)-(1)(x)'/(x)^@

OpenStudy (callisto):

You can use power for d/dx (1/x) \[\frac{d}{dx} \frac{1}{x} = \frac{d}{dx} (x^{-1}) = (-1)x^{-1-1} =...?\]

OpenStudy (anonymous):

-x^(-2)?

OpenStudy (callisto):

Yes.

OpenStudy (anonymous):

ok so it will be e^(-x^(-2))?

OpenStudy (callisto):

The power of e would not change when you differentiate e^(something)!

OpenStudy (callisto):

\[\frac{dy}{dx} = \frac{dy}{du}\times \frac{du}{dx} = \frac{d}{du}e^u \times \frac{d}{dx}(\frac{1}{x})= ...?\] u = 1/x and you found d/dx(1/x). So......

OpenStudy (anonymous):

i'm lost with this one

OpenStudy (callisto):

Where?

OpenStudy (anonymous):

are you asking for the derivative of (1/x) above? i dont understand what exactly you are asking

OpenStudy (callisto):

I asked derivative of 1/x because you didn't get that right. To get the derivative of derivative of (x^(2))(e^(1/x)), it should be like this: \[y = x^2e^{\frac{1}{x}}\]\[y' = x^2\frac{d}{dx}e^{\frac{1}{x}} + e^{\frac{1}{x}}\frac{d}{dx}(x^2)\] Then, then you need to work out what d/dx (e^(1/x)) is and d/dx (x^2) are. The later one is easy, and you got that right. The problem is to find d/dx (e^(1/x)) So, let u = 1/x \[\frac{d}{dx}e^{\frac{1}{x}} = \frac{d}{du}(e^u) \times \frac{d}{dx} (\frac{1}{x}) = ...?\]

OpenStudy (callisto):

Is that clear? Do you understand what we are working on?

OpenStudy (anonymous):

but isn't derivative of (1/x) = -x^(-2) ? i thought we got that right

OpenStudy (anonymous):

wait. give me second

OpenStudy (callisto):

Yes. d/dx (1/x) = -x^(-2) = -1/x^2

OpenStudy (anonymous):

so you are asking derivative of e^(1/x) * -1/x^2?

OpenStudy (callisto):

No. Instead, derivative of e^(1/x), which is equal to e^(1/x) * -1/x^2. Do you understand how to get it?

OpenStudy (anonymous):

so looking at the original question question my answer should be (2x)(e^1/x)-(x^2)(e^(1/x))?

OpenStudy (anonymous):

derivative of (x^(2))(e^(1/x)) is the original question

OpenStudy (callisto):

No... The second part is not correct. For the second part, you need to find the derivative of e^(1/x). What is it?

OpenStudy (anonymous):

isn't it e^(1/x)

OpenStudy (anonymous):

e^(1/x) * -1/x^2?

OpenStudy (callisto):

It is *NOT* e^(1/x) e^(1/x) * -1/x^2 ^How do you get it?

OpenStudy (anonymous):

because derivative of 1/x = -1/x^2

OpenStudy (callisto):

Yup.. So, can you solve \(y' = x^2\frac{d}{dx}e^{\frac{1}{x}} + e^{\frac{1}{x}}\frac{d}{dx}(x^2)\) now?

OpenStudy (callisto):

^That is your derivative btw.

OpenStudy (anonymous):

yes i had already gotten the first part. just the second was the problem

OpenStudy (callisto):

Just show us what you've got.

OpenStudy (anonymous):

final answer for the original question is (2x)(e^(1/x))-(x^2)(1/x^2)(e^(1/x))

OpenStudy (anonymous):

this is what i got

OpenStudy (callisto):

That looks pretty cool~ But you can simplify the last term.

OpenStudy (anonymous):

i don't think we have to. but you can show me if you want

OpenStudy (callisto):

\[(2x)(e^{\frac{1}{x}})-(x^2)\frac{1}{x^2}(e^{\frac{1}{x}}) = (2x)(e^{\frac{1}{x}})-\frac{x^2}{x^2}(e^{\frac{1}{x}}) =...?\]

OpenStudy (anonymous):

x^2 / x^2 gets cancelled out = 1

OpenStudy (callisto):

Yes.

OpenStudy (anonymous):

perfect.

OpenStudy (anonymous):

Thank you very much for your help

OpenStudy (anonymous):

you explain concepts very well

OpenStudy (callisto):

You're welcome. I hope I didn't confuse you :S \[(2x)(e^{\frac{1}{x}})-\frac{x^2}{x^2}(e^{\frac{1}{x}}) =(2x)(e^{\frac{1}{x}})-e^{\frac{1}{x}}=e^{\frac{1}{x}}(2x-1)\]That looks nice :)

OpenStudy (anonymous):

No, not confusing at all

OpenStudy (anonymous):

yes that looks better

OpenStudy (callisto):

:)